### v1

# X-LeBench: A Benchmark for Extremely Long Egocentric Video Understanding

Abstract
Long-form egocentric video understanding provides rich contextual information and unique insights into long-term human behaviors, holding significant potential for applications in embodied intelligence, long-term activity analysis, and personalized assistive technologies. However, existing benchmark datasets primarily focus on single, short-duration videos or moderately long videos up to dozens of minutes, leaving a substantial gap in evaluating extensive, ultra-long egocentric video recordings. To address this, we introduce X-LeBench, a novel benchmark dataset specifically crafted for evaluating tasks on extremely long egocentric video recordings. Leveraging the advanced text processing capabilities of large language models (LLMs), X-LeBench develops a life-logging simulation pipeline that produces realistic, coherent daily plans aligned with real-world video data. This approach enables the flexible integration of synthetic daily plans with real-world footage from Ego4D—a massive-scale egocentric video dataset covers a wide range of daily life scenarios—resulting in 432 simulated video life logs that mirror realistic daily activities in contextually rich scenarios. The video life-log durations span from 23 minutes to 16.4 hours. The evaluation of several baseline systems and multimodal large language models (MLLMs) reveals their poor performance across the board, highlighting the inherent challenges of long-form egocentric video understanding and underscoring the need for more advanced models.
1 Introduction
Understanding long-form egocentric videos captured from a first-person perspective over extended periods holds significant potential for advancing various domains such as embodied intelligence, long-term activity analysis, and personalized assistive technologies [30, 25, 29]. These videos provide rich contextual information and unique insights into human behaviors as they unfold naturally throughout the day. The ability to analyze such extensive recordings is key to developing more personalized agent systems that can construct long-term memory from users’ daily recordings, better understand human actions, anticipate needs, and interact more seamlessly in real-world settings [24, 17, 31].
However, most existing datasets [37, 5, 18, 20, 43, 27], predominantly focus on short individual video clips captured from third-person views, failing to document detailed individual activities that make them insufficient for in-depth human-centered research. Egocentric video understanding has recently gained more attention, with several notable benchmarks and datasets being proposed [8, 9, 12, 50, 26, 13, 25, 10]. Ego4D [12] features massive-scale egocentric videos covering a wide range of daily activities, with video durations ranging from 5 seconds to 7 hours. Several subsequent research [26, 13, 16, 35, 6] based on Ego4D have also emerged. They fall short of capturing the full scope of long-term daily human activities, as they mainly focus on isolated clips or recordings. This limitation hinders the development and evaluation of models designed to process ultra-long video streams. In particular, it restricts the ability to challenge current long-form video understanding systems and models that construct long-term memory from video recordings and retrieve relevant information in response to user queries. Without extensive, continuous contextual data, evaluating the robustness of their performance becomes difficult.
Creating benchmark datasets that span several hours of egocentric video is necessary but presents significant challenges. Data acquisition is a primary hurdle, recording dozens or even hundreds of hours-long egocentric videos is labor-intensive and raises privacy concerns, as it requires participants to wear recording devices for extended periods while going about their daily lives. Device limitations add to the complexity, as continuous video capture demands substantial storage capacity and consistent performance from recording equipment, which can affect data quality and reliability. Additionally, ensuring annotation quality in long-form videos is daunting. Manually annotating extensive recordings is time-consuming and can lead to annotator fatigue, compromising the detail and accuracy of the labels needed for training and evaluating complex models.
To address these challenges, we introduce X-LeBench, a versatile and scalable benchmark dataset specifically designed for evaluating tasks on extremely long egocentric videos. X-LeBench features a life-logging simulation pipeline that synthesizes new data from existing short (a few seconds) or moderately long (a few hours) videos. It employs large language models (LLMs) to dynamically generate realistic daily plans aligned with real-world activities based on adjustable input settings. Specifically, this simulation integrates synthetic daily plans with actual footage from Ego4D, matches them in three core information aspects, and then dynamically and iteratively optimizes the simulation process based on the retrieved information, resulting in simulated video life logs that mirror daily activities in rich contexts with their duration extended to dozens of hours. A generated sample of the video life-log simulation is shown in Fig. 1, which consists of video recordings at different timestamps. Notably, X-LeBench offers a customizable and scalable design that allows to synthesis data of different lengths and content, facilitating future research of tailored needs.
Figure 1: Example of generated video life logs in X-LeBench. Generated video life logs consist of multiple videos with corresponding timestamps. The visualization shows the time organization and content allocation of data.
Our initial evaluations of baseline systems and multimodal LLM (MLLM) show consistently poor performance on X-LeBench. These results highlight the inherent difficulties of long-form egocentric video understanding and underscore the need for more advanced models. X-LeBench aims to facilitate research in this area by providing a comprehensive benchmark, encouraging the development of solutions that can effectively interpret and analyze ultra-long egocentric videos.
Our contributions are:
  • We present X-LeBench, the first benchmark dataset that encompasses ultra-long egocentric video recordings.
  • We introduce a novel and customizable pipeline that simulates realistic, hours-long egocentric videos by integrating synthetic daily plans with real-world footage.
  • We conduct extensive evaluations of existing models on X-LeBench, revealing significant performance gaps and highlighting the challenges that future research must address.
2 Related Works
Table 1: The comparison of various benchmarks.
Dataset Avg. Duration (mins) Annotation Scheme Egocentric
MVBench [19] 0.27 Auto No
ActivityNet-QA [5] 1.85 Manual No
EgoSchema [26] 3 Auto&Manual Yes
EgoPlan-Bench2 [33] up to 5 Auto&Manual Yes
MovieChat-1K [36] 9.4 Manual No
MLVU [49] 12 Auto&Manual Partial
Video-MME (Short) [11] 1.37 Manual No
Video-MME (Medium) 9.38
Video-MME (Long) 39.76
HourVideo [6] 45.7 Auto&Manual Yes
InfiniBench [2] 52.59 Auto&Manual No
LVBench [42] 68.35 Manual No
Ours (Short) 142 Auto&Manual Yes
Ours (Medium) 319
Ours (Long) 516
2.1 Egocentric Video Understanding Benchmarks
Egocentric video understanding has received increasing attention in embodied intelligence, long-term activity analysis, and personalized assistive technologies [7, 14]. Ego4D [12] addresses the demand for extensive egocentric data by providing over 3,670 hours of egocentric videos across 74 locations. It enables significant advancements, but presents videos in isolated segments, lacking coherent, continuous recordings under real-life contexts.
Related works [4, 23, 31, 16] extend Ego4D, further exploring various applications of egocentric video understanding. EgoSchema [26] offers an egocentric video question-answering benchmark with over 5,000 curated multiple-choice pairs but focuses on 3-minute clips, insufficient for the evaluation of extended temporal understanding.
The Aria Everyday Activities dataset [25] offers multimodal egocentric data,
HourVideo [6] curates videos of dozens of minutes from Ego4D for evaluating hour-long video-language understanding. EgoPlan-Bench2 [33] assesses planning capabilities of MLLMs. However, these works either lack coherent, continuous daily-life recordings over long time spans or face data scale and diversity limitations. In contrast, our work addresses these gaps by building an ultra-long, continuous daily-life egocentric video dataset.
2.2 Long-form Video Understanding Benchmarks
The definition of “long” in video understanding varies across benchmarks, ranging from a few minutes to several hours, with most focusing on third-person videos. As shown in Tab. 1, EgoSchema [26] defines 3-minute videos as long using their proposed certificate length, Video-MME [11] classifies videos with a length of 30-60 minutes as long. HourVideo [6] and LVBench [42] specifically propose benchmarks for long video understanding, and define videos longer than 20 and 30 minutes as long videos respectively. MLVU [49] offers videos of diversified lengths with a 12-minute average length, providing comprehensive evaluation tasks for MLLMs’ long video understanding capabilities. InfiniBench [2] pushes the boundary with 50-minute videos and over 108,000 question-answer pairs, posing significant challenges for leading AI models.
Despite recent advances, these datasets are neither sufficiently long for evaluating ultra-long video processing nor focused on egocentric content, leaving a gap in addressing this emerging real-world application. Our work addresses this by redefining “long video” to include multi-hour, contextually consistent egocentric recordings with sufficient annotations to advance model and system development.
2.3 LLM-assisted Annotation Scheme
Traditional annotation processes require extensive human effort, but advances in LLMs and MLLMs [1, 44, 3, 38, 40, 21] have enabled their use in benchmarks. As shown in Tab. 1, a range of video understanding benchmarks leverage LLMs/MLLMs to automate the annotation process [26, 34, 32]. For instance, EgoSchema [26] generates question-answer pairs by querying an LLM with human-annotated narrations and tailored prompts, MVBench [19] uses LLMs to generate annotations for diverse video datasets.
In our work, the video duration extends to more than ten hours, making manual annotation virtually impractical and prone to fatigue-related errors. To address this, we also adopt an automatic and manual annotation scheme. First, we adapt Ego4D’s manual annotations to fit our task design. Then, we use LLMs to consolidate summaries of 5-minute video clips from Ego4D at various granularities, creating single-video, multi-video, and holistic-level summaries (detailed in Sec. 3.2.2).
3 X-LeBench
In real-world scenarios, life-logging with egocentric devices is constrained by hardware limitations and privacy concerns, preventing continuous recording. Thus, obtaining long-term life-logging video records is inevitably a combination of multiple video segments captured at different times. Building on this insight, we develop the life-logging simulation pipeline. By leveraging the text processing capabilities of LLMs and the extensive annotations of the Ego4D dataset, we generate simulated video life logs that realistically reflect daily activities and maintain contextual coherence, eventually culminate in the creation of our X-LeBench, detailed in the following sections.
Figure 2: Overview of life-logging simulation pipeline. Stage 1: Generation of personalized persona profiles and daily plan chunks based on predefined parameters. Stage 2: Core information (time, scene, content) extraction of videos in the video selection library. Stage 3: Matching and retrieval of daily plan chunks with videos, and life-log simulations are iteratively refined through reflection, resulting in the final optimized output.
3.1 Life-logging Simulation Pipeline
As shown in Fig. 2, We constructed the pipeline by implementing the following three stages:
Stage 1 - Persona Generation. We dynamically generate personalized character profiles based on the input of different locations and the myers-briggs type indicators (MBTI) [28], generating basic background information including the character’s personality traits, lifestyle, hobbies and general daily routines. This information is the initial character’s settings to facilitate subsequent diverse and unique generations of detailed daily plans. We then generate a detailed daily activities agenda that matches the initial settings, and further use it to generate the daily plan chunks, where each chunk includes the activity plan of the character in different time periods. During this stage, we use GPT-4o [15] for character generation, setting 9 different locations based on the demographics of Ego4D dataset, and 16 MBTI types for each location, resulting in a total of 144 different persona profiles.
Stage 2 - Video Information Extraction. The Ego4d dataset provides video content that spans hundreds of scenarios, allowing a glimpse into daily life around the world, and is labeled with dense video summaries for every 5-minute video clip. We construct our video selection library by carefully selecting 7852 videos from the dataset that are labeled with scenarios and clip-level summaries and not redacted. Then we extract three core aspects of video information from the library, including the time, scene, and the main activities presented in these videos. Specifically, we use Gemini 1.5 Pro [39] to analyze videos. To facilitate simulation with reasonable environment settings and lighting conditions, videos are categorized into different scene environments: indoor, outdoor, and mixed, as well as different time periods: daytime, nighttime, twilight, and uncertain. It is worth noting that the main scenes of some videos are indoors, making it difficult to recognize their time period information, for this kind of videos, we assign the special label “uncertain”. As for the extraction of content information, we directly extract the scenario information (i.e. main activity description) from Ego4D as the overview of the video content, and use Gemini to consolidate a series of 5-minute clip-level summaries into a 1-sentence summary of each video.
Figure 3: Dataset statistics. (a): Video selection library consists of 7852 videos from Ego4D dataset, covering 135 scenarios, here shows the top 30. (b): X-LeBench has video life-log durations span from 23 minutes to 16.4 hours, here shows distribution of duration range across different data categories. (c) and (d) show the distribution of scene and time information categories of videos in video selection library, respectively. (e): Statistical information of duration lengths of different data categories, including minimum, mean, median and maximum values. (f): The distribution of number of videos with different occurrences.
Stage 3 - New Dataset Generation. In this stage, we generate the new dataset by linking the generated daily plan chunks (from Stage 1) with the real video information (from Stage 2). For each chunk, we retrieve a video that aligns with its time, scene, and content, constructing the video life-log simulation. The selection process follows this rule:
$V={\underset{V\in\mathbf{M_{v}}}{{\arg\max}}\ S(\mathbf{M_{v}}(C_{t},C_{scene},C_{scenario}),C_{desc})}$ (1)
where $V$ is the selected video for chunk $C$, determined by the alignment with $C$’s core information, which is time ($C_{t}$), scene ($C_{scene}$), scenario ($C_{scenario}$), and 1-sentence summary description ($C_{desc}$). $\mathbf{M_{v}(\cdot)}$ denotes the operation of matching videos based on the provided chunk information. $S(\cdot)$ represents the sentence similarity computation between $C_{desc}$ and 1-sentence summary set of matching videos. Specifically, $C_{t}$ and $C_{scene}$ are inferred from chunk timestamps and content, while $C_{scenario}$ is inferred using the dataset’s predefined scenario list (i.e. the dataset reference in Fig. 2).
To efficiently identify the most suitable videos from the large-scale video selection library, we first apply a coarse-grained filtering step, i.e., the operation $\mathbf{M_{v}}(C_{t},C_{scene},C_{scenario})$, a rule-based matching to narrow down the candidate video set. Finally, to refine the selection, we calculate the sentence similarity between the 1-sentence summary of $C$ and the summaries in the candidate video set ($\mathbf{M_{v}}$), ensuring alignment across time, scene, and content dimensions.
In addition, to enhance the contextual reasonableness and coherence of overall video life logs, we preferentially select videos recorded within the same country for each simulation. Furthermore, an iterative optimizing strategy is adopted, i.e., after each matching of a daily plan chunk, we update the life-log simulation memory with the information of selected video, which acts as the real daily activities record of the persona. This information, along with the remaining chunks to be simulated, is provided to the LLM. The LLM uses this context to reasonably adjust and update the subsequent daily plan chunks, ensuring that the overall simulation maintains logical coherence and contextual alignment.
Statistics. Our life-logging simulation pipeline integrates flexible parameterization settings. By specifying the location and MBTI-type information, we can get persona profiles of different styles, resulting in diverse simulations. Also, we can customize different nnumber of chuncks to retrieve and get video life logs of various video compositions. In this work, we set three different chunk numbers for simulation, namely 4, 9, and 15 (corresponding to short, medium, and long life-log categories), for each location-MBTI combination generated persona, thus obtaining 432 extremely long video life logs.
As summarized in Tab. 2, Fig. 3 (b) and (e), the duration statistics across categories show significant diversity: short data are mostly 1-2 hours, medium data are mostly 4-5 hours, and long data are mostly 6-8 hours, with the average duration of 2.37, 5.32 and 8.6 hours respectively.
Table 2: Statistics of dataset duration length and quality evaluation score.
Life-log Category Max (min) Mean (min) Min (min)
Short 580 142 23
Medium 760 319 118
Long 984 516 220
Our constructed video selection library consisting of 7852 videos, covers a wide range of daily life scenarios and scenes across different times of the day, as shown in Fig. 3 (a), (c) and (d), a total of 135 scenarios are covered, here we show the top 30 most frequent occurrences in Fig. 3 (a). The statistics of the time period and scene information are shown in Fig. 3 (c), and (d) respectively. In addition, due to the commonality of human activities (e.g., eating, cooking), certain videos in the video selection library are selected multiple times. As shown in Fig. 3 (f), the vast majority of videos in X-LeBench appear less than 5 times, with a small number of videos appearing more than 5 times, reflecting the diversity of our simulations while preserving the representativeness of repeated contexts.
Table 3: The number of annotations across different tasks.
Object-related Retrieval
People-related
Retrieval
Action Counting Summary Ordering
1444 583 5295 4032
Moment Retrieval Summarization
single-video multi-video holistic
9869 4032 1245 432
3.2 Benchmark Tasks
The goal is to comprehensively evaluate the extremely long video understanding capabilities of current systems and models. As shown in Tab. 4, X-LeBench introduces a suite of evaluation tasks specifically designed for daily-life long-form egocentric videos. These tasks encompass object-, people-, and moment-related temporal localization, multi-level summarization, action counting, and summary ordering.
Temporal Localization
Object-related Retrieval Q: What did I put on the table in the record provided for the 21:56 - 22:25 time period? A: [22:01:55 - 22:02:03], [22:02:20 - 22:02:42], …
People-related Retrieval Q: Who did I talk to in the living room in the record provided for the 21:56 - 22:25 time period? A: [22:02:35 - 22:02:39], [22:03:30 - 22:03:50], …
[…]
Summarization
Single-video Summarization Q: Summarize the activities performed in the recordings provided for the time period 22:13 - 22:25. A: C interacts with others and writes in a house and a studying room.
Multi-video Summarization Q: Summarize the activities performed in the recordings provided in the morning (before 12:00). A: C engages in various activities, including preparing meals, riding…
Holistic Summarization Q: Summarize the activities performed in all provided life-log recordings. A: C engages in a variety of activities throughout the day, including
      preparing meals, riding, visiting restaurants, …
Counting
Action Counting Q: In the record provided for the 07:25 - 07:48 time period, how many times have each of the A: wipe soap: 2; remove cheese: 1; …
      actions in the following list been performed in the 0-28 second record of the period?
      Action list: 1. wipe soap, 2. remove cheese, …
Ordering
Summary ordering Q: Please rank the following summaries of camera wearer C’s activities in order of presentation of A: Correct order of the summaries:
      the life-log recordings.       7, 2, 11, 5, 8, 1, 6, 3, 4, 0, 12, 9, 13, 10, 14.
      Summary 0: C climbs, interacts with others, and observes climbing activities in various indoor…
      Summary 14: C interacts with others and writes in a house and a studying room.
Table 4: Tasks in this benchmark and their corresponding examples. Q and A denote query and answer examples, respectively.
The full-length video is presented to the system only once before querying. The system is required to extract features or frames from the extremely long video and store them in a buffer, responding to query based solely on the stored information. This approach presents a significant challenge to current video understanding systems, which are typically designed to process short videos and thus cannot store information in the presented videos in buffers for future use. Moreover, in real-world scenarios, corresponding answers often recur across multiple time segments as the recording time span increases, so we explicitly annotate the queried time period for each task. This ensures more precise, unambiguous answers, and provides a solution to address the limitations and incompleteness of annotations in existing single-video datasets.
After generating the life-log footage, we selected the corresponding annotations from the Ego4D dataset [12] and adapted them for our novel tasks. Detailed descriptions of each task are provided below. A summary of the number of annotations compiled for each task is presented in Tab. 3.
3.2.1 Temporal Localization
Task Design and Annotation: Temporal localization is critical for enabling virtual personal assistants to handle user queries effectively in egocentric videos. This task, akin to Episodic Memory [12] and Natural Language Video Localization [47, 46], is uniquely challenging for ultra-long videos due to their extensive duration and diverse content. We divide this task into three subtasks based on the type of information being retrieved: objects, people, and moments. For object- and people-related retrieval, we adopt the annotations from the Episodic Memory task (specifically, the Natural Language Query subtask) in Ego4D. Specifically, we retrieve the annotations corresponding to the videos selected by the daily plan chunks (as generated in Sec. 3.1), classify them into topics based on the query templates defined in Ego4D, and remapped the timestamps to align with the virtual time used in the daily plan.
Evaluation Metrics: Following previous work on temporal localization [12, 47, 46], we employed a top-$x$ recall with an intersection over union threshold (“recall@x, tIoU”). $x$ and $t$ are predefined variables that indicate the number of retrieved results we look at and the IoU threshold, respectively. Here, we set $x$ to 5 and $t$ to 0.3.
3.2.2 Summarization
Task Design and Annotation: Summarization is a fundamental task in numerous benchmarks [42, 49, 2], In our context, it provides insights at varying granularities—specific chunks, predefined periods (e.g., morning, afternoon, evening), and holistic summaries. These diverse scopes challenge systems to summarize at multiple summary levels in ultra-long videos. We divide the task into three granularity levels: single-video, multi-video, and holistic. Single-video level focuses on summarizing individual chunks from the daily plan. Multi-video level involves summarizing three predefined time periods: morning, afternoon, and evening. [^1] Morning: before 12 p.m., afternoon: 12 p.m. to 5 p.m., evening: after 5 p.m. Holistic level requires an overall summary of the entire daily footage. To generate summaries at these levels, we employed a hierarchical approach inspired by [16]. Single-video summaries are initially consolidated during the earlier stages of processing (as described in  Sec. 3.1). Using these, we generate multi-video summaries by querying a LLM. The resulting multi-video summaries were then aggregated to produce the holistic summary in a similar manner.
Evaluation Metrics: Although there is an increasing trend to use LLM-based scoring systems for the summarization task (such as in [2]), we found such approaches to be highly unstable and sometimes produce inconsistent scores. Therefore, we instead use traditional summarization metrics such as Rouge-L [22] and the BERT score [48].
3.2.3 Counting
Task Design and Annotation: Counting task tests a system’s ability to accurately retrieve, interpret, and aggregate fine-grained information from multiple parts of the inputs. This task involves counting the occurrences of specific actions within the inputs to reflect user needs for recalling how many times a particular action was performed. Future extensions could involve counting objects or people appearing in the footage.
Evaluation Metrics: We employed a simple accuracy metric. Accuracy was calculated as the ratio of correctly counted actions to the total number of actions queried.
3.2.4 Ordering
Task Design and Annotation: Ordering task evaluates a model’s capability to recognize temporal relationships between main activities across video records in ultra-long video life logs. Unlike prior studies focusing on fine-grained actions, this task introduces the novel challenge of ordering events based on single-video summaries. Models are tasked with ordering shuffled single-video summaries of entire life logs, posing significant challenges to the system’s temporal understanding ability in long-context inputs. This setup ensures the initial sequence of events is obscured, requiring the model to reconstruct the correct temporal order.
Evaluation Metrics: To evaluate this ordering task, we propose to use a simple accuracy metric. This accuracy is the ratio between the number of correctly predicted order and the total number of summaries.
4 Experiments
4.1 Settings
X-LeBench includes various task types, as outlined in Sec. 3.2, encompassing four categories: temporal localization, summarization, counting, and ordering. These tasks are further divided into eight subtasks, with their respective input-output formats detailed in Tab. 4. Given the interleaved, ultra-long context with chronological attributes in our dataset (Fig. 1), we uniformly construct time-stamped prompts as the long-context input following the format shown in Fig. 4.
Figure 4: Example of input context.
To better align with real-world usage, our tasks are defined as requiring free-form textual outputs rather than close-set answers for multiple-choice questions. This approach increases practical usability while significantly raising the complexity of tasks, as it demands more nuanced understanding and reasoning. To ensure the evaluability of outputs, the prompts provide extended context along with task-specific instructions and output format requirements for each task type, standardizing the expected responses. For the temporal localization and summarization tasks, each query is independently evaluated, while for the counting and ordering tasks, we aggregate the query contents of each data entry to action or order lists for a unified assessment. Notably, the computational cost and time associated with analyzing ultra-long contexts, along with their overall poor performance and limited analytical value, we test only a limited set of temporal localization tasks.
Overall, X-LeBench presents significant challenges for current video understanding systems, which often lack robust temporal reasoning or retrieval capabilities and are typically constrained to short contexts or video understanding. In this section, we evaluated X-LeBench using three multimodal approaches:
(1) Gemini-1.5-Flash [39]: This native multimodal model excels in handling ultra-long contexts and is trained jointly on multimodal data. It is suitable to directly evaluate X-LeBench in an end-to-end way. In our experiments, the model is set to a temperature of 0.1, and 1200 frames of each data are uniformly sampled as input context.
(2) Socratic Models [45]: Most advanced multimodal systems cannot process ultra-long videos. Inspired by related works [6, 45], we divide videos into 30-second segments and generate captions for these segments using the Qwen-VL-7B model [41], captioning segments at 1 frame per second and resolution settings of $560\times 420$. The captions are then aggregated with timestamps as life-log records, forming the input for executing long-context understanding tasks. The Gemini-1.5-Flash model is used for the long-context question-answering tasks on these textual life log records, with the same parameters as in method (1).
(3) Retrieve-Socratic: Considering the inherent temporal nature of our dataset, we propose an enhanced version of the Socratic Models approach. This method incorporates a rule-based filtering mechanism that extracts only the time segments relevant to the task, thereby excluding irrelevant temporal information from the life-log records. By narrowing the contextual focus, this approach aims to improve efficiency and relevance in long-context understanding. The extracting contextual information is then processed using the same Socratic Models’ pipeline and parameter settings as in method (2).
Method Data Category Object-related Retrieval (%) People-related Retrieval (%) Moment Retrieval (%) Summarization (RougeL / BertScore) (100) AVG. Temporal Localization (%) AVG. Summarization (100) Action Counting (%) Ordering (%)
single multi-video holistic
Gemini-1.5-Flash Short 1.67 0.00 10.26 25.02 / 88.16 20.44 / 86.81 19.96 / 85.65 4.20 23.47 / 87.53 13.87 93.70
Medium 1.92 6.90 10.53 26.00 / 88.55 17.69 / 85.64 16.10 / 83.37 5.88 24.15 / 87.75 16.74 44.66
Long 1.75 11.11 8.33 26.24 / 88.67 17.46 / 84.83 14.12 / 82.96 6.45 24.72 / 87.98 16.76 24.40
Socratic Models Short 8.33 25.00 17.95 24.12 / 88.21 17.73 / 86.09 22.65 / 87.92 14.28 23.01 / 87.87 9.53 85.59
Medium 5.77 20.69 18.42 24.66 / 88.24 15.66 / 85.34 23.79 / 88.53 13.45 23.51 / 87.92 14.04 55.63
Long 12.28 22.22 10.83 24.53 / 88.13 15.70 / 85.28 21.94 / 87.80 11.83 23.39 / 87.79 19.85 23.33
Retrieve - Socratic Short 8.33 25.00 17.95 23.06 / 88.11 20.60 / 86.78 22.97 / 88.05 14.28 22.74 / 87.93 9.92 87.67
Medium 9.62 17.24 7.89 22.20 / 87.85 18.16 / 86.12 24.18 / 88.67 10.92 21.91 / 87.72 9.38 54.55
Long 10.53 66.67 14.17 22.33 / 87.87 18.36 / 86.28 22.94 / 88.31 15.59 21.98 / 87.74 9.77 24.17
Gemini-1.5-Flash All 1.77 5.17 9.14 25.97 / 88.55 18.28 / 85.60 16.87 / 84.06 5.66 24.30 / 87.82 16.44 41.96
Socratic Models All 8.87 22.41 13.70 24.51 / 88.18 16.07 / 85.45 22.79 / 88.08 12.97 23.36 / 87.86 17.15 42.61
Retrieve - Socratic All 9.47 27.59 13.71 22.39 / 87.90 18.72 / 86.32 23.36 / 88.34 13.92 22.08 / 87.77 9.67 43.01
Table 5: Evaluation results on X-LeBench. Including temporal localization tasks (object-related, people-related and moment retrieval), summarization tasks (single-video, multi-video and holistic level), action counting and summary ordering task. AVG. temporal localization is the average performance of all temporal localization tasks, AVG. summarization is the average performance of all summarization tasks.
Figure 5: Performance comparison on X-LeBench. (a): The performance comparison of each method across tasks for different data categories. (b): The performance comparison across different data categories for different tasks.
4.2 Results
As shown in Fig. 5 and Tab. 5, we report the performance of 3 methods on the “short”, “medium”, “long” and “all” categories of data across different tasks. The “all” category represents an aggregated evaluation of the former 3 data types, offering a more comprehensive and holistic assessment. The evaluation metrics for each task are detailed in Sec. 3.2.
Overall Performance. As shown in last three rows of Tab. 5 and Fig. 5 (a), the Retrieve-Socratic method demonstrates superior performance in most tasks. Notably, in the temporal localization task, the Retrieve-Socratic method achieves an average recall that surpasses Gemini-1.5-Flash by 8.26%. In summarization tasks, Gemini-1.5-Flash exhibits stronger performance, specifically, for the single-video summarization subtask, it outperforms other methods. However, due to input token constraints, longer videos require larger sampling intervals, leading to increasing information loss. As a result, Gemini-1.5-Flash’s performance is notably inferior to that of the Socratic Models and Retrieve-Socratic methods for temporal localization tasks, where temporal information of finer granularity is crucial. In holistic summarization and ordering tasks, the queried time periods are global. Under such conditions, the Socratic Models and Retrieve-Socratic method yield nearly identical results, demonstrating strong consistency in performance across these methods. Interestingly, our analysis of the specific results indicates that the Retrieve-Socratic method, which employs temporal information filtering to reduce contextual information, tends to output 0 more frequently for the counting task. Consequently, its performance on the counting task is notably worse than other methods. This suggests that excessive pruning of contextual information may inadvertently lead to unpredictable result bias.
Impact of Data Types. As illustrated in Fig. 5 (b) and Tab. 5, the performance trends across data types are not uniform. These trends may be influenced by varying task-specific objectives and understanding burden imposed by different data types on each method. Overall, both Socratic and Retrieve-Socratic demonstrate more stable performance trends across data types compared to Gemini-1.5-Flash. This suggests that the processing capability of textual information is more robust when dealing with such long-context information. For the ordering task, all methods achieve strong performance on short videos (accuracy exceeding 85%). However, as video length increases, performance across all methods declines significantly, with accuracy dropping below 25% for long data. For all subtasks, Gemini-1.5-Flash’s performance on long data is worse than on short or medium data, except on people-related retrieval, single-video summarization and counting tasks, Socratic Models and Retrieve-Socratic have the same trends as Gemini-1.5-Flash, except on object-related retrieval and counting tasks, and people- and object-related retrieval tasks, respectively.
Remarks. Notably, due to the information security constraints, Gemini-1.5-Flash refuses to respond to certain data. The refusal rates for short, medium, and long data are 20.14%, 27.08%, and 30.56%, respectively. In contrast, for the Socratic Models and Retrieve-Socratic methods, where video data is first converted into textual information before being processed by Gemini, the refusal rate is reduced to 0%. Additionally, in the multi-video summarization task, a significant portion of responses are invalid even when timestamps are included in input context, and the query time is explicitly specified in the question (e.g., “12:00 to 17:00”). Typical invalid responses include statements like “There is no information available within the specified time period”. We classify such responses as failed replies, which are also excluded from evaluation. The failure rates of the multi-video summarization task are 55.28%, 51.89%, and 57.03% for Gemini-1.5-Flash, Socratic Models and Retrieve-Socratic, respectively.
Summary of Key Findings.
  • Performance on Long Videos: All methods face significant challenges with long video inputs, underscoring the difficulty of handling ultra-long temporal contexts.
  • Retrieve-Socratic Method: Excels in temporal localization tasks, demonstrating strong fine-grained temporal understanding. However, excessive information pruning introduces biases in counting tasks.
  • Gemini-1.5-Flash: Shows superior summarization performance but struggles with temporal localization under long-context inputs due to token limitations and information loss.
  • Performance Across Data Lengths: Performance trends vary across input lengths, with Socratic Models and Retrieve-Socratic exhibiting greater robustness compared to Gemini-1.5-Flash. Suggesting that, under ultra-long video inputs with temporal information, converting information to text is a more efficient and robust way for long-context understanding.
Overall, these results highlight the significant challenges posed by long video input, also emphasize the need for improved cross-length generalization and enhanced temporal information handling.
5 Conclusion
We present a customizable life-logging simulation pipeline for generating realistic and contextually coherent simulated video life logs. Introducing X-LeBench, the first benchmark for ultra-long egocentric video tasks, we curate 432 video life logs from Ego4D, categorized into short, medium, and long durations. X-LeBench also includes a suite of evaluation tasks encompassing temporal localization, summarization, counting, and ordering. By providing this challenging benchmark, we hope to promote the development of advanced models and schemes capable of understanding long-form egocentric videos and tackling the complexities of temporal information.
Despite our efforts to construct an extremely long video dataset with contextually coherent activity contents, limitations remain due to the scarcity and insufficient diversity of available data. For details of videos, contextual inconsistencies exist within the dataset, such as different kitchen or office settings within the same data. However, since our focus is on long-term activity content, these discrepancies are beyond the scope of this work. Future work could involve expanding both the modalities and time spans of collected data through diverse devices and sources to enhance a comprehensive understanding of long-form human activity.